(d) The table below shows the marginal utility schedule for orange (X)      and apples (Y). Suppose that oranges and apples are the only two  commodities available, the price of orange is Shs 10 , the price of   apple            is Shs 10, and individual income is Shs 800 per time period and all is spent.

Quantity

Marginal utility of X

Marginal Utility of Y

10

20

30

40

50

60

70

110

100

90

80

70

60

50

190

170

150

130

120

100

80

 

           Required:

           

(i)    How many units of good X and good Y will this utility maximizing consumer buy if the level of income is Shs 14                                                                                  (3 marks)



1
Expert's answer
2021-12-02T10:38:08-0500

(i) The utility is maximized when:

"MUx\/Px = MUy\/Py" and

"Px\u00d7X + Py\u00d7Y = I."

Both equalities are true, when 20 units of X and 60 units of Y are consumed, because 100/10 = 100/10 and 10×20 + 10×60 = 800.